# Local asymptotics of selection models with applications in Bayesian selective inference

### v2

@@ 5 Discussion
To date, Gaussianity assumptions have formed the basis of most analysis of inferential properties of selection models and the techniques of selective inference. The asymptotic analysis presented in this paper establishes formal connections between the behavior of selective inference methodologies in more general parametric settings and of their analogues in the corresponding Gaussian limits. The main focus for future work will be to utilize these connections to examine operational and theoretical optimality properties of procedures advocated for inference in the presence of selection. In particular, it will be of methodological importance to explore extensions of the main result to regression settings.
@@ A.1 Proof of Theorem 1
Consider the sequence of random vectors $[Z,U^{*}_{n}]$, where $Z\sim N(h,I_{\theta}^{-1})$ and $U^{*}_{n}\mid Z\sim\text{Bernoulli}\{p^{*}_{n}(Z;\theta)\}$, so that $Z\mid U_{n}^{*}=1$ has selection function $p^{*}_{n}(z;\theta)$. The analogous functions for this model are given by
@@
(A) For all $K>0$, the collection $\{\exp(h^{T}Z_{n})\colon\|h\|\leq K,n\in\mathbb{N}\}$ is uniformly integrable under $\theta$. To see this, consider the moment generating function $M(x)=E_{\theta}[\exp\{x^{T}\triangledown l_{\theta}(Y_{1})\}]$. We have $E_{\theta}[\exp(h^{T}Z_{n})]=M(n^{-1/2}h)^{n}$, so it suffices to prove that
$\sup\{E_{\theta}[\exp(h^{T}Z_{n})^{2}]=M(n^{-1/2}2h)^{n}\colon|h|\leq K,n\in\mathbb{N}\}<\infty.$ (24)
@@
$\sup_{\|h\|\leq K,n\in\mathbb{N}}M\left(\frac{2h}{n^{1/2}}\right)^{n}\leq\exp\left\{2K^{2}\sup_{\|x\|\leq K^{\prime}}\triangledown^{2}\log M(x)\right\}<\infty.$ (26)
Clearly, $\{\exp(h^{T}I_{\theta}Z_{n})\colon\|h\|\leq K,n\in\mathbb{N}\}$ is also uniformly integrable for all $K>0$. Thus, for every $\varepsilon>0$, there exists a $C_{1}>0$ such that
@@
$\sup_{\|h\|\leq K,n\in\mathbb{N}}E\left[\exp(h^{T}I_{\theta}Z)\mathbf{1}\{\exp(h^{T}I_{\theta}Z)>C_{2}\}\right]<\varepsilon.$ (28)
@@
which satisfies $\|g_{n}(\cdot;h)\|_{\infty}\leq C\|p^{*}_{n}(\cdot;h)\|_{\infty}$. If the score has an absolutely continuous component, by Theorem 2.5 of Bally and Caramellino (2016), $d_{\text{TV}}(Z_{n},Z)\to 0$, so $E_{\theta}[g_{n}(Z;h)]-E_{\theta}[g_{n}(Z_{n};h)]=o(1)$. Otherwise, define $A(h)=\{z\colon\exp(h^{T}I_{\theta}z)\leq C\}$. We have that
@@
where we have used that $\|fg\|_{\text{L}}\leq\|f\|_{\infty}\|g\|_{\text{L}}+\|g\|_{\infty}\|f\|_{\text{L}}$. Since the bounded-Lipschitz distance metrizes convergence in distribution, we also get $E_{\theta}[g_{n}(Z;h)]-E_{\theta}[g_{n}(Z_{n};h)]=o(1)$. Thus, for all $\varepsilon>0$,
$\left|E_{\theta}[\exp(h_{n}Z_{n})p^{*}_{n}(Z_{n};\theta)]-E_{0}[\exp(h_{n}Z)p^{*}_{n}(Z;\theta)]\right|\leq 2\varepsilon\|p^{*}_{n}(\cdot;\theta)\|_{\infty}+o(1),$ (33)
@@
Putting all together, using $\dot{=}$ to indicate equality up to an $o(1)$ term,
@@ A.4 Proof of Proposition 2
For notational convenience, we are going to consider the two-dimensional case with
$\Sigma=\begin{pmatrix}\sigma_{1}^{2}&\rho\sigma_{1}\sigma_{2}\\ \rho\sigma_{1}\sigma_{2}&\sigma_{2}^{2}\end{pmatrix},$
and parameter of interest $\theta_{1}$, though all the steps are readily generalizable to $p>2$. First, let us transform the data to
$Z=\begin{pmatrix}1&0\\ -c\lambda&c\end{pmatrix}Y\sim N\left(\begin{pmatrix}\theta_{1}\\ c[\theta_{2}-\lambda\theta_{1}]\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right)\equiv N\left(\begin{pmatrix}\psi\\ \chi\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right),$
where
$\lambda$ $=$ $\rho\frac{\sigma_{2}}{\sigma_{1}},$
$c$ $=$ $\left(\sigma_{2}^{2}\left[1+\frac{\rho^{2}}{\sigma_{1}^{2}}(1-2\sigma_{1}^{2})\right]\right)^{-1/2},$
and $\psi=\theta_{1}$ and $\chi=c(\theta_{2}-\lambda\theta_{1})$ are the parameter of interest and the nuisance parameters, respectively. Let $\tilde{\varphi}(\psi,\chi)$ be the selection probability in the $(\psi,\chi)$ parametrization. We want to show that $\Pi(\psi_{0}\mid Z)$ is not distributed as $U(0,1)$ under any value of $\psi_{0}$.
The selective posterior density for $\psi$, for a given prior $\pi(\psi,\chi)$, has
$\pi(\psi\mid z_{1},z_{2})\propto\phi(\psi-z_{1})g(\psi;z_{2}),\quad g(\psi;z_{2})=\int_{-\infty}^{\infty}\pi(\psi,\chi)\frac{\phi(\chi-z_{2})}{\tilde{\varphi}(\psi,\chi)}\mathrm{d}\chi.$
We are going to show first that $\Pi(\psi_{0}\mid Z_{1},Z_{2})\mid Z_{2}=z_{2}\not\sim U(0,1)$ for some $z_{2}$ when $\psi_{0}$ is the true value of the interest parameter, irrespective of the value of $\psi_{0}$. The selective density of $Z_{1}$ given $Z_{2}=z_{2}$ is
$f_{\psi}(z_{1}\mid z_{2},u)=\frac{\phi(\psi-z_{1})\tilde{p}(z_{1},z_{2})}{\tilde{\varphi}(\psi;z_{2})},\quad\tilde{\varphi}(\psi;z_{2})=\int_{-\infty}^{\infty}\phi(\psi-z_{1})\tilde{p}(z_{1},z_{2})\mathrm{d}z_{1},$
where $\tilde{p}(z_{1},z_{2})$ is the selection function for the transformed data $Z$. Lindley (1958) showed that, for a one-dimensional model to admit an exact probability-matching prior, there needs to exist a change of variables $X=x(Z_{1})$ and a reparametrization $\tau=\tau(\psi)$ such that the model for $X\mid\tau$ is a location model. However, no transformation of $\psi$ can turn the conditional model of $Z_{1}\mid Z_{2}=z_{2}$ into a location model unless $\tilde{\varphi}(\psi;z_{2})\propto 1$, as the likelihood is not expressible in the form $L\{\tau-x^{-1}(X)\}$ for any $L$. Furthermore, since $\tilde{\varphi}(\psi,\chi)=E_{\chi}[\tilde{\varphi}(\psi;Z_{2})]$, we have, by the assumption on the selection probability, that
$E_{\chi}\left[\frac{\partial}{\partial\psi}\tilde{\varphi}(\psi;Z_{2})\right]$ $=$ $\frac{\partial}{\partial\psi}\tilde{\varphi}(\psi,\chi)$
$=$ $\frac{\partial}{\partial\theta_{1}}\varphi(\theta_{1},\theta_{2})\frac{\partial\theta_{1}}{\partial\psi}+\frac{\partial}{\partial\theta_{2}}\varphi(\theta_{1},\theta_{2})\frac{\partial\theta_{2}}{\partial\psi}$
$=$ $\frac{\partial}{\partial\theta_{1}}\varphi(\theta_{1},\theta_{2})+\lambda\frac{\partial}{\partial\theta_{2}}\varphi(\theta_{1},\theta_{2})$
$=$ $\frac{1}{\sigma_{1}^{2}}\left[\text{Cov}(Y_{1},Y_{1})\frac{\partial}{\partial\theta_{1}}\varphi(\theta_{1},\theta_{2})+\text{Cov}(Y_{1},Y_{2})\frac{\partial}{\partial\theta_{2}}\varphi(\theta_{1},\theta_{2})\right]$
$\neq$ 00 0
for all $(\psi,\chi)$. It follows that $\psi\to\tilde{\varphi}(\psi;Z_{2})$ cannot be constant in $\psi$ with probability one. If $z_{2}$ is such that $\tilde{\varphi}(\psi;Z_{2})$ is not constant, there does not exist any prior $\pi(\psi;z_{2})$ that provides exact frequentist matching when attached to the conditional likelihood of $Z_{1}\mid Z_{2}=z_{2}$. Since $\pi(\psi\mid z_{1},z_{2})$ can be obtained by combining the conditional likelihood with $\pi(\psi;z_{2})\propto\tilde{\varphi}(\psi;z_{2})g(\psi;z_{2})$, it follows that $\Pi(\psi_{0}\mid Z_{1},Z_{2})$ is not be uniform conditionally on $Z_{2}=z_{2}$, for a set of $z_{2}$’s of non-zero probability.
Finally, from the latter observation we show that $\Pi(\psi_{0}\mid Z_{1},Z_{2})$ cannot be uniform under all possible values of the nuisance parameter $\chi_{0}$, thus failing to satisfy the required matching condition. If $\Pi(\psi_{0}\mid Z_{1},Z_{2})\sim U(0,1)$, then
$P_{\psi_{0},\chi_{0}}\{\Pi(\psi_{0}\mid Z_{1},Z_{2})\leq\alpha\}=E_{\psi_{0},\chi_{0}}\left[P_{\psi_{0}}\{\Pi(\psi_{0}\mid Z_{1},Z_{2})\leq\alpha\mid Z_{2}\}\right]=\alpha,\quad\forall\alpha\in(0,1).$
However, $Z_{2}$ is complete for $\chi$, as it is the sufficient statistic of a natural exponential family. Hence, this equality can only hold if $P_{\psi_{0}}\{\Pi(\psi_{0}\mid Z_{1},Z_{2})\leq\alpha\mid Z_{2}\}=\alpha$ for all $\alpha$ with probability one, which contradicts the earlier statement.
@@ A.5 Proof of Proposition 3
Assume for simplicity that $\sigma^{2}=1$. As in the proof of Theorem 1, consider the model for $(Y,U)$ where $U\mid Y\sim\text{Bernoulli}\{p(Y)\}$, and let $H(\theta;y)=P_{\theta}(Y\geq y\mid U=1)$. Note that $H(\theta;Y)\sim U(0,1)$ under $\theta$. Rewrite this as
@@
The second factor is also increasing. Assume without loss that $\|p\|_{\infty}=1$ (as the model is invariant to rescaling) and note that the selection condition can be restated as $Y\geq T$, where $T$ is independent of $Y$ and $P(T\leq t)=p(t)$. Write

### v3

@@ 5 Discussion
To date, Gaussianity assumptions have formed the basis of most analysis of inferential properties of selection models and the techniques of selective inference. The asymptotic analysis presented in this paper establishes formal connections between the behavior of selective inference methodologies in more general parametric settings and of their analogues in the corresponding Gaussian limits. An important focus for future work will be to utilize these connections to examine operational and theoretical optimality properties of procedures advocated for inference in the presence of selection. In particular, it will be of methodological importance to explore extensions of the main result to regression settings.
@@
Another important avenue for future research concerns the extension of the present results to selection regimes with vanishing selection probabilities. This would provide a unified theoretical basis for the study selection models with IID data, and would substantially broaden the applicability of the framework in high-dimensional contexts.
6 Acknowledgments
We would like to thank the Associate Editor and three anonymous referees for their valuable comments and suggestions on a previous version, which have improved the clarity and quality of the work.
@@ A.1 Proof of Theorem 1
Consider the sequence of random vectors $Z,U^{*}_{n}]$, where $Z\sim N(h,I_{\theta}^{-1})$ and $U^{*}_{n}\mid Z\sim\text{Bernoulli}\{p^{*}_{n}(Z;\theta)\}$, so that $Z\mid U_{n}^{*}=1$ has selection function $p^{*}_{n}(z;\theta)$. The analogous functions for this model are given by
@@
(A) For all $K>0$, the collection $\{\exp(h^{T}I_{\theta}Z_{n})\colon\|h\|\leq K,n\in\mathbb{N}\}$ is uniformly integrable under $\theta$. To see this, consider the moment generating function $M(x)=E_{\theta}[\exp\{x^{T}\triangledown l_{\theta}(Y_{1})\}]$. We have $E_{\theta}[\exp(h^{T}I_{\theta}Z_{n})]=M(n^{-1/2}h)^{n}$, so it suffices to prove that
$\sup\{E_{\theta}[\exp(h^{T}Z_{n})^{2}]=M(n^{-1/2}2h)^{n}\colon\|h\|\leq K,n\in\mathbb{N}\}<\infty.$ (28)
@@
$\sup_{\|h\|\leq K,n\in\mathbb{N}}M\left(\frac{2h}{n^{1/2}}\right)^{n}\leq\exp\left\{2K^{2}\sup_{\|x\|\leq K^{\prime}}\|\triangledown^{2}\log M(x)\|_{\text{op}}\right\}<\infty,$ (30)
where $\|\cdot\|_{\text{op}}$ is the operator norm. Thus, for every $\varepsilon>0$, there exists a $C_{1}>0$ such that
@@
$\sup_{\|h\|\leq K}E_{0}\left[\exp(h^{T}I_{\theta}Z)\mathbf{1}\{\exp(h^{T}I_{\theta}Z)>C_{2}\}\right]<\varepsilon.$ (32)
@@
which satisfies $\|g_{n}(\cdot;h)\|_{\infty}\leq C\|p^{*}_{n}(\cdot;h)\|_{\infty}$. If the score has an absolutely continuous component, by Theorem 2.5 of Bally and Caramellino (2016), $d_{\text{TV}}(Z_{n},Z)\to 0$, so $E_{\theta}[g_{n}(Z_{n};h)]-E_{0}[g_{n}(Z;h)]=o(1)$. Otherwise, define $A(h)=\{z\colon\exp(h^{T}I_{\theta}z)\leq C\}$. We have that
@@
where we have used that $\|fg\|_{\text{L}}\leq\|f\|_{\infty}\|g\|_{\text{L}}+\|g\|_{\infty}\|f\|_{\text{L}}$. Since the bounded-Lipschitz distance metrizes convergence in distribution, we also get $E_{\theta}[g_{n}(Z_{n};h)]-E_{0}[g_{n}(Z;h)]=o(1)$. Thus, for all $\varepsilon>0$,
$\left|E_{\theta}[\exp(h_{n}I_{\theta}Z_{n})p^{*}_{n}(Z_{n};\theta)]-E_{0}[\exp(h_{n}I_{\theta}Z)p^{*}_{n}(Z;\theta)]\right|\leq 2\varepsilon\|p^{*}_{n}(\cdot;\theta)\|_{\infty}+o(1),$ (37)
@@
[…]
@@ A.4 Proof of Proposition 4
If a prior is probability-matching for $\theta$, it is also probability-matching for every subparameter of interest $\theta_{i}$. Without loss of generality, let the parameter of interest be $\theta_{1}$, $\pi(\theta)$ be a valid arbitrary prior density, not necessarily proper, and denote by $\Pi(\theta_{1}\mid y)$ the corresponding marginal posterior distribution function. The prior is probability-matching for $\theta_{1}$ if $\Pi(\theta_{1}\mid Y)\sim U(0,1)$ under all true parameters $\theta\in\mathbb{R}^{p}$. We are going to assume $\Sigma=I_{p}$, which is analytically easier, and we will consider the general case at the end.
First, we show that $\Pi(\theta_{1}\mid Y)\sim U(0,1)$ for all $\theta\in\mathbb{R}^{p}\iff\Pi(\theta_{1}\mid Y)\mid Y_{-1}\sim U(0,1)$ almost surely for all $\theta\in\mathbb{R}^{p}$. To see this, write
$P_{\theta}\{\Pi(\theta_{1}\mid Y)\leq\alpha\}-\alpha=E_{\theta}[P_{\theta_{1}}\{\Pi(\theta_{1}\mid Y)\leq\alpha\mid Y_{-1}\}-\alpha]=0.$
Since $Y_{-1}$ is complete for $\theta_{-1}$ for every fixed value of $\theta_{1}$ (it is the sufficient statistic of an exponential family with natural parameter $\theta_{-1}$), it follows that
$P_{\theta_{1}}\{\Pi(\theta_{1}\mid Y)\leq\alpha\mid Y_{-1}\}=\alpha$
with probability 1 for all $\theta$.
Now, define $H(\theta_{1};y)=P_{\theta_{1}}(Y_{1}\geq y_{1}\mid y_{-1})$. We are going to show that the previous condition can only hold if $\Pi(\theta_{1}\mid Y)=H(\theta_{1};Y)$ almost surely for all $\theta_{1}\in\mathbb{R}$. Clearly, $H(\theta_{1};Y)\mid y_{-1}\sim U(0,1)$ for all values of $\theta$ and for all values of $y_{-1}$ for which the conditional distribution is well-defined, and $H(\theta_{1};y)$ is a valid distribution function for $\theta_{1}$ for every fixed $y$, because the conditional model on $Y_{1}\mid y_{-1}$ is stochastically increasing in $\theta_{1}$. Additionally, we know that $\Pi(\theta_{1}\mid y)$ is strictly decreasing in $y_{1}$ regardless of the prior, as the posterior density can be expressed in the form
$\pi(\theta_{1}\mid y)=\frac{\phi(y_{1}-\theta_{1})g(\theta_{1};y_{-1})}{h(y)},\quad g(\theta_{1};y_{-1})=\int_{-\infty}^{\infty}\frac{\pi(\theta)}{\varphi(\theta)}f_{\theta_{-1}}(y_{-1})\mathrm{d}\theta_{-1}.$
Let $l_{\alpha}(\theta,y_{-1})$ be defined by
$P_{\theta_{1}}\{\Pi(\theta_{1}\mid Y)\leq\alpha\mid y_{-1}\}=P_{\theta_{1}}\{Y_{1}\geq l_{\alpha}(\theta,y_{-1})\mid y_{-1}\}=H(\theta_{1};l_{\alpha}(\theta,y_{-1}),y_{-1})=\alpha.$
For $\alpha=\Pi(\theta_{1}\mid y)$ we obtain, from the last equality, $H(\theta_{1};y)=\Pi(\theta_{1}\mid y)$.
Differentiating this equality with respect to $\theta_{1}$ and $y_{1}$, we get
$\frac{\partial^{2}}{\partial\theta_{1}\partial y_{1}}H(\theta_{1};y)$ $=$ $\frac{\partial}{\partial y_{1}}\frac{\phi(y_{1}-\theta_{1})g(\theta_{1};y_{-1})}{h(y)}$
$-\frac{\partial}{\partial\theta_{1}}\frac{\phi(y_{1}-\theta_{1})p(y)}{\varphi(\theta_{1};y_{-1})}$ $=$ $\frac{\partial}{\partial y_{1}}\frac{\phi(y_{1}-\theta_{1})g(\theta_{1};y_{-1})}{h(y)},$
where $\varphi(\theta_{1};y_{-1})=\int_{\mathbb{R}}\phi(y_{1}-\theta_{1})p(y)\mathrm{d}y_{1}$, so
$\frac{p(y)}{\varphi(\theta_{1};y_{-1})}\left[\theta_{1}-y_{1}+\frac{\partial}{\partial\theta_{1}}\log\varphi(\theta_{1};y_{-1})\right]=\ \frac{g(\theta_{1};y_{-1})}{h(y)}\left[\theta_{1}-y_{1}-\frac{\partial}{\partial y_{1}}\log h(y)\right].$
For a fixed $y_{-1}$, rewrite this equation as
$a(\theta_{1})\left[\theta_{1}-y_{1}+\frac{\partial}{\partial\theta_{1}}\log\varphi(\theta_{1};y_{-1})\right]=b(y_{1})\left[\theta_{1}-y_{1}-\frac{\partial}{\partial y_{1}}\log h(y)\right],$
where $a(\theta_{1})^{-1}=\varphi(\theta_{1};y_{-1})g(\theta_{1};y_{-1})$ and $b(y_{1})^{-1}=h(y)p(y)$. Differentiating w.r.t. $\theta_{1}$ we obtain
$a^{\prime}(\theta_{1})\left[\theta_{1}-y_{1}+\frac{\partial}{\partial\theta_{1}}\log\varphi(\theta_{1};y_{-1})\right]+a(\theta_{1})\left[1+\frac{\partial^{2}}{\partial\theta_{1}^{2}}\log\varphi(\theta_{1};y_{-1})\right]=b(y_{1})>0.$
The left hand side is linear in $y_{1}$, with slope $a^{\prime}(\theta_{1})=b^{\prime}(y_{1})$. This implies that $a(\theta_{1})$ is linear, but since it is also strictly positive for all $\theta_{1}$, the only possibility is that $a^{\prime}(\theta)=b^{\prime}(y_{1})=0$ for all $\theta_{1}$. Thus,
$1+\frac{\partial^{2}}{\partial\theta_{1}^{2}}\log\varphi(\theta_{1};y_{-1})\propto 1\iff\frac{\partial^{3}}{\partial\theta_{1}^{3}}\log\varphi(\theta_{1};y_{-1})=0$
for all $\theta_{1}$ and $y_{-1}$ such that $\varphi(\theta_{1};y_{-1})>0$. By the invertibility of the Weierstrass transform, this condition can be written directly in terms of the selection function, which needs to be almost everywhere quadratic in $y_{1}$ of the form
$\log p(y)=a(y_{-1})y_{1}^{2}+b(y_{-1})y_{1}+c(y_{-1}).$
Applying the same argument to any orthonormal transformation $\tilde{Y}=MY\sim N(\tilde{\theta},I_{p})$, with selection function $p(M^{T}y)$, it follows that $\log p(y)$ is quadratic in every possible direction, so
$\log p(y)=y^{T}Ay+b^{T}y+c\quad\text{for some }A\in\mathbb{R}^{p\times p},b\in\mathbb{R}^{p},c\in\mathbb{R}.$
To conclude, consider a general covariance matrix $\Sigma$ and define $Z=\Sigma^{-1/2}Y\sim N(\tilde{\theta},I_{p})$, with selection function $\tilde{p}(z)=p(\Sigma^{1/2}y)$. Probability-matching is invariant to the parametrization, and for the latter model to admit such a prior,
$\log p(y)=\log\tilde{p}(\Sigma^{-1/2}z)=y^{T}\Sigma^{-1/2}A\Sigma^{-1/2}y+b^{T}\Sigma^{-1/2}+c$
for some $A\in\mathbb{R}^{p\times p},b\in\mathbb{R}^{p},c\in\mathbb{R}$, so the quadratic form holds for $p(y)$ as well.
Since the selection function depends on $A$ via $y^{T}Ay$, only the symmetric part of $A$ is relevant, so we can assume it is symmetric without loss of generality. The further assumption that $\Sigma^{-1}-2A\succ 0$ is required for the selective model,
$Y\sim N\left((\Sigma^{-1}-2A)^{-1}(\Sigma^{-1}\theta+b),(\Sigma^{-1}-2A)^{-1}\right),$
to be well-defined.
@@ A.5 Proof of Proposition 5
Assume for simplicity that $\sigma^{2}=1$, and assume without loss that $\|p\|_{\infty}=1$ (as the model is invariant to rescaling). As in the proof of Theorem 1, consider the model for $(Y,U)$ where $U\mid Y\sim\text{Bernoulli}\{p(Y)\}$, and let $H(\theta;y)=P_{\theta}(Y\geq y\mid U=1)$. Note that $H(\theta;Y)\sim U(0,1)$ under $\theta$. Rewrite this as
@@
The second factor is also increasing. Since the selection function is right-continuous and $\|p\|_{\infty}=1$, note that the selection condition can be restated as $Y\geq T$, where $T$ is independent of $Y$ and $P(T\leq t)=p(t)$. Write